### v2

# Hyperparameter Importance Analysis for Multi-Objective AutoML

Abstract
Hyperparameter optimization plays a pivotal role in enhancing the predictive performance and generalization capabilities of ML models. However, in many applications, we do not only care about predictive performance but also about objectives such as inference time, memory, or energy consumption. In such MOO scenarios, determining the importance of hyperparameters poses a significant challenge due to the complex interplay between the conflicting objectives. In this paper, we propose the first method for assessing the importance of hyperparameters in the context of multi-objective hyperparameter optimization. Our approach leverages surrogate-based hyperparameter importance (HPI) measures, i.e. fANOVA and ablation paths, to provide insights into the impact of hyperparameters on the optimization objectives. Specifically, we compute the a-priori scalarization of the objectives and determine the importance of the hyperparameters for different objective tradeoffs. Through extensive empirical evaluations on diverse benchmark datasets with three different objectives paired with accuracy, namely time, demographic parity, and energy consumption, we demonstrate the effectiveness and robustness of our proposed method. Our findings not only offer valuable guidance for hyperparameter tuning in MOO tasks but also contribute to advancing the understanding of HPI in complex optimization scenarios.
[…]
2239
1 Introduction
The selection of appropriate hyperparameter configurations significantly impacts a model’s ability to capture underlying patterns in the data and produce accurate predictions. Optimizing hyperparameters is one of the main focus areas of Automated Machine Learning (AutoML) [21, 7]. Given a use case and a model, it is usually not known which hyperparameters are worth tuning to achieve a good performance. Providing insights into this (e.g., [36]) is valuable since it allows the design of better configuration spaces and gives a better understanding of the learning dynamics of machine learnirng (ML) algorithms. Hyperparameter Importance (HPI) offers a systematic method to gain insights into the influence of hyperparameters on the model’s performance. Understanding and optimizing hyperparameters are crucial steps in building effective ML models. They allow us to fine-tune our algorithms, improve performance, and achieve better generalization.
So far, most efforts in hyperparameter optimization (HPO) have been focused on single-objective optimization, mainly targeting the predictive performance of models. Recently, there has been a trend towards multi-objective HPO [17, 3, 13, 15, 6, 22], which allows the consideration of several objectives, including fairness, memory consumption, training time, inference time, and energy consumption. In a multi-objective optimization (MOO) scenario, where we optimize conflicting objectives simultaneously, it becomes not only more challenging to approximate the unknown Pareto front of non-dominated configurations but also to discern the relative significance of individual hyperparameters.
Figure 1: Overview of the MO-fANOVA method. On the left, the importance of each hyperparameter for Objective 1 is shown. On the right, our extension for the importance of each hyperparameter for different weightings of the objectives is displayed exemplarily.
Conventional methods for assessing HPI rely on univariate sensitivity analysis techniques, like variance-based methods [20] or partial dependence plots [28]. For instance, techniques like fANOVA have been widely used to decompose the variance of model performance into contributions from individual hyperparameters [20]. While effective for single-objective optimization, these methods may not capture the intricate interactions among multiple objectives in multi-objective settings. So far, there are no methods that assess the importance of hyperparameters within the context of MOO.
Figure 2: Overview of the MO-ablation path analysis. On the left, an exemplary Pareto front is displayed, with several ablation paths going from the default configuration to different configurations on the Pareto front. Every path is associated with a weighting of the objectives and thus gives a different value for the difference in performance per hyperparameter. We convert this to the plot on the right, where the total performance for different weightings is displayed as a stacked plot of hyperparameter contributions.
In this paper, we propose a novel method called weighted multi-objective hyperparameter importance (MO-HPI). Although our approach is in principle usable with any surrogate-based HPI measure, we specifically focus on two widely-used methods, fANOVA [20] and ablation path analysis [14, 4]. Our methodology involves training surrogate models, e.g., random forests, with the hyperparameter configuration data and the respective objective results. We use weighted sums of the objectives as target variables, allowing us to effectively capture the change in importance from one objective to another. Figure 1 shows the concept of MO-fANOVA and Figure 2 gives an overview of the idea for the MO-ablation path analysis.
This work aims to propose a method to analyze MOO results and demonstrate that it is reasonable in the context of ML. To reach that goal, we pose the following research question: How to convert surrogate-based HPI methods to meaningfully measure the importance of hyperparameters within multi-objective optimization? We further frame the following subquestions: Are the results intuitively correct? How do the proposed methods compare? Can we gain new insights into the impact of hyperparameters on the optimization process that previous methods could not provide?
To answer our research questions and validate the efficacy of our proposed method, we conduct three empirical evaluations on the well-known benchmark datasets MNIST [9], Adult Census Income [2] and CIFAR10 [24]. To show that the approach can be applied to diverse objectives, we consider three objectives paired with accuracy: training time, demographic parity, and energy consumption, and discuss the results for each of them accordingly.
By advancing the understanding of HPI in complex optimization scenarios, our method offers a valuable technique that can aid researchers and practitioners in tackling challenging MOO problems more effectively.
2 Related Work
Several techniques are available for evaluating the importance of hyperparameters. Surrogate models, such as Gaussian process models [31] or random forests [8], have been employed to approximate the relationship between hyperparameters and performance. They can predict the performance of a given hyperparameter configuration based on an empirical dataset of configurations and their performances. Breiman [8] introduced how random forests can attribute importance by observing performance changes when removing attributes. Forward selection [19] uses this idea by selecting the hyperparameters that most affect the surrogate’s performance by starting from no hyperparameters and iteratively adding the most impactful ones. Local Parameter Importance [5] studies the performance changes of a configuration along each hyperparameter by considering the variance in performance when changing the hyperparameter. Recent works [1, 32] use Shapley values [35] to measure the HPI in Bayesian optimization inspired by the fact that Shapley values can quantify the hyperparameter attributions for the acquisition function.
Another method is ablation path analysis [14, 4], which compares the default configuration and the optimized configuration to measure hyperparameter contributions. It creates an ablation path from the default configuration to a target configuration by changing the hyperparameter with the largest increase in performance in each iteration. We note that these ablation paths do not follow traditional ablation studies in which only a single hyperparameter is changed and all others are fixed; but this allows a full path through the configuration space from the default configuration to a target configuration. While there is only one path in a single-objective scenario, there are potentially many in the MOO setting (cf. Figure 2).
fANOVA [20] identifies the importance of individual hyperparameters and interactions among them. For each hyperparameter, it measures how it contributes to the variance in performance. This is done, e.g., by training a random forest as a surrogate model and subsequently decomposing the variance of each tree into contributions to each subset of hyperparameters. Based on these methods, several works have been published. For example, van Rijn and Hutter [36] applied HPI across datasets. PED-ANOVA [38] improved fANOVA using Pearson divergence to work better on arbitrary subspaces of the search space, e.g. the subspace of the top-performing configurations. This paper contributes to this body of work by extending the methodology to MOO.
3 Weighted HPI for Multiple Objectives
Our main idea is simple and intriguing given the insight that we are interested in the HPI for different objectives tradeoffs. The main challenge lies in (i) how to map this to different HPI methods, (ii) how to compute this efficiently, (iii) how to visualize this, and (iv) how to interpret this. Challenges (i)-(iii) are described in this section and (iv) follows as part of the following experiment and discussion sections. Overall, our work employs a framework for assessing the influence of hyperparameters on multiple objectives using any surrogate-based HPI analysis methods, allowing it to be efficient in computing analyses. Given a performance dataset obtained by an MO-HPO optimizer, surrogate models are trained to calculate the HPI. In the following, the steps will be explained in more detail. W.l.o.g. we discuss our approaches for two competing objectives; the approach can be straightforwardly extended to more objectives.
3.1 Data Preparation.
The performance dataset should contain model configurations with their performance on several objectives. First, each of the objectives is normalized using the min-max normalization. Next, the configuration data is converted to numerical values to deal with categorical data and nan values.
3.2 Weighting Scheme
The objectives are weighted by the tradeoffs of the optimized configurations. Those are the configurations where the objective values are on the Pareto front. The configurations on the Pareto front are the set of non-dominated solutions. To retrieve the corresponding weighting, only Pareto-efficient objective pairs $(o_{1},o_{2})$ of the non-dominated solutions are considered. Subsequently, the normalized objective pairs are scaled to sum up to 1 by adding the pair up and dividing each value by the total, to be used as weighting.
3.3 Multi-Objective fANOVA
Algorithm 1 Multi-Objective fANOVA
1:$(\vec{O_{1}},\vec{O_{2}})\leftarrow$ normalized objective values
2:$X\leftarrow$ encoded configuration data
3:$\vec{W}\leftarrow$ is_pareto_efficient($(\vec{O_{1}},\vec{O_{2}})$)
4:for each $\vec{w}$ in $\vec{W}$ do
5:     $Y_{w}\leftarrow w_{1}\cdot\vec{O_{1}}+w_{2}\cdot\vec{O_{2}}$
6:     $\mathcal{S}\leftarrow$ surrogate model trained with $(X,Y_{w})$
7:     Calculate fANOVA importances with $\mathcal{S}$ following [20]
8:end for
A single target variable (Y) is created for each weighting $(w_{1},w_{2})$ by summing the normalized weighted objectives values $(\vec{O_{1}},\vec{O_{2}})$:
$Y_{w}=w_{1}\cdot\vec{O_{1}}+w_{2}\cdot\vec{O_{2}}$ (1)
Given the weighting and the prepared data, MO-fANOVA can be calculated. Algorithm 1 shows the procedure in detail. A probabilistic surrogate model is trained for each weighting. The model takes the encoded configurations as input and the weighted normalized sum of the objectives as the target variable. After that, the fANOVA importance is calculated using the trained model for the corresponding weighting. See Figure 1 for an exemplary depiction of the outcome.
3.4 Multi-Objective Ablation Path Analysis
As shown by Biedenkapp et al. [4], ablation paths, i.e., the path of flipping the value of a hyperparameter from a given default configuration to a target configuration, can be efficiently approximated by using a surrogate model trained on meta-data collected by algorithm configuration or HPO. However, the standard procedure of computing these ablation paths by a greedy scheme [14] only considers a single objective.
Algorithm 2 Multi-Objective Ablation Path Analysis
Input: evaluated configurations $\lambda\in\vec{\Lambda}$, corresponding evaluated objective values $(\vec{O_{1}},\vec{O_{2}})$, hyperparameters $hps$
1:$(\vec{O_{1}},\vec{O_{2}})\leftarrow$ objective values
2:$\vec{\Lambda}\leftarrow$ encoded configuration data
3:$\vec{W}\leftarrow$ is_pareto_efficient($(\vec{O_{1}},\vec{O_{2}})$)
4:for each $obj$ in $(\vec{O_{1}},\vec{O_{2}})$ do $\triangleright$ Not normalized
5:     $\vec{\mathcal{S}}[{obj}]\leftarrow$ random forest trained with $\vec{\Lambda},Y_{\text{obj}}$
6:end for
7:for each $\vec{w}$ in $\vec{W}$ do
8:     $\lambda_{\text{best}}\in\underset{\vec{\Lambda}}{\operatorname*{arg\,min}}(w_{1}\cdot\vec{O_{1}}+w_{2}\cdot\vec{O_{2}})$
9:     $r_{\text{previous}}\leftarrow\sum(\vec{w}\times\vec{\mathcal{S}}(\lambda_{\text{default}}))$
10:     $r_{\text{min}}\leftarrow r_{\text{previous}}$
11:     while $r_{\text{min}}\leq r_{\text{previous}}$ do $\triangleright$ As long as we find improvements
12:         for $hp$ in $hps$ do
13:              $\lambda_{\text{default}}[hp]\leftarrow\lambda_{\text{best}}[hp]$
14:              for each $\mathcal{S}_{\text{obj}}$ in $\vec{\mathcal{S}}$ do
15:                  $\vec{r}[obj]\leftarrow$ normalized $\mathcal{S}_{\text{obj}}(\lambda_{\text{default}})$
16:              end for
17:              $r_{\text{total}}\leftarrow\sum(\vec{w}\times\vec{r})$triangleright Scalarized performance
18:              if $r_{\text{total}}<r_{\text{min}}$ then
19:                  $hp_{\text{min}}\leftarrow hp$triangleright Most important hyperparameter
20:                  $r_{\text{min}}\leftarrow r_{\text{total}}$
21:              end if
22:              reset $\lambda_{\text{default}}$
23:         end for
24:         Add $hp_{\text{min}}$ with $r_{\text{min}}$ to ablation path
25:         Remove $hp_{\text{min}}$ from $hps$
26:     end while
27:end for
We propose to extend the ablation path analysis for multiple objectives as follows, see also Algorithm 2. For both objectives, a surrogate model is trained with the encoded configuration data and the objectives before normalization (line 4). For each weighting, the ablation path is calculated (starting from line 7). The incumbent $\lambda_{best}$ is the corresponding configuration to the minimum of the scalarized target variable $Y_{w}$ of Equation 1 (line 8). Starting from the default configuration, each hyperparameter value is changed individually to the value of the incumbent configuration, and the performance is estimated with the trained surrogate models (lines 12-16). The results of the models are normalized with the same normalization as the respective objective values, weighted, and summed up as performance (line 15 and 17). The hyperparameter value leading to the highest difference in performance (toward the optimization objective) will be changed in the configuration (lines 18-21 and line 24). The ablation is repeated with the altered configuration and the remaining hyperparameters until the incumbent configuration is reached (line 11). The difference in performance is recorded for each weighting.
4 Experiments
Using our MO-HPI approach, we evaluated three HPO problems with different data, models, and objectives. The overview can be seen in Table 1. We chose rather simple benchmarks and models so that the results of the MO-HPI approach are easier to validate by common knowledge.
Table 1: Overview of the experiments
Name Dataset Model Objective 1 Objective 2
time MNIST MLP 1-Accuracy Training Time
fairness Adult Census MLP 1-Accuracy Demographic Parity
energy CIFAR10 ResNet 1-Accuracy Energy Consumption
4.1 General Setup
The basic setup was the same for all experiments. We used SMAC [18, 26] as one of the state-of-the-art HPO tools [12] to perform HPO. The multi-objective algorithm ParEGO [23] with Random Online Aggressive Racing (ROAR) [18] and 1000 trials was used for optimization. ROAR randomly selects a configuration from the hyperparameter space and only keeps track of the top 10 non-dominated configurations. We note that our primary goal was not to achieve the best possible multi-objective performance but to use a sufficient amount of performance data generated by an HPO tool. Since Moosbauer et al [28] showed post-hoc analysis of HPO runs can be biased by too strong exploitation, we focus here on unbiased randomly generated data. For both HPI methods, we trained random forests with 100 trees as surrogate models.
4.2 Experiment: Time
Figure 3: Results for the time experiment. The Pareto front is on the left (error vs. training time in seconds) with the red dot being the default performance. The MO-fANOVA results are in the middle and the MO-ablation path analysis is on the right. The x-axis corresponds to the weighting of the minimum error objective.
Table 2: Hyperparameter Space for Multi-Layer Perceptron
Hyperparameter Range Scale Default
n_layer 1-5 linear 3
n_neurons (same
for all layers)
8-256 log 132
activation logistic, tanh, relu - tanh
initial_lr 0.0001-0.1 log 0.01
alpha 0.0001-1.0 log 0.1
beta_1 0.1-1.0 log 0.5
beta_2 0.1-1.0 log 0.5
epsilon 1e-10-1e-06 log 1e-8
The first experiment used the MNIST dataset with Multi-layer Perceptron (MLP) classifiers based on Sci-kit learn’s implementation [30]. The objective “training time” was measured by the wallclock time the classifier used for training. The error was measured by subtracting the accuracy on the test set from 1. The MLPs were configured with 50 maximum epochs and Adam as the DL optimizer. SMAC tuned the MLPs hyperparameter w.r.t. the configuration space in Table 2. Figure 3 shows the Pareto front (left), the MO-fANOVA (middle) and the MO-ablation path analysis (right) results for the time experiment. Note that for the x-axis only the weighting for the error objective is displayed. Since the total weighting always sums to 1, the opposite weighting applies to the respective other objective — in this case, training time (so $w_{2}$ = $1-w_{1}$). The ablation path analysis plot is stacked, with each segment representing the performance increase attributed to a specific hyperparameter for each weighting. The blue part shows the performance of the default configuration.
4.3 Experiment: Fairness
Figure 4: Results for the fairness experiment. The Pareto front is on the left (error vs. demographic parity) with the red dot being the default performance. The MO-fANOVA results is in the middle and the MO-ablation path analysis is on the right. The x-axis corresponds to the weighting of the minimum error objective.
The second experiment used the Adults Census Income dataset [2] which contains several variables about US citizens with the binary target variable annual income higher or lower than 50’000\$. We only used the numeric and binary nominal variables of the dataset. It contains the sensitive variables “sex” and “race”. We calculated the fairness loss as the second objective for the experiment based on the prediction of the “race” variable using demographic parity (DP). It is calculated by the absolute difference of the mean proportions of positive predictions $y$ in each group, where the groups are defined by a sensitive variable $s$.
$DP=\left|\frac{\sum_{i=1}^{n}y_{{}_{i}}(s_{i}=0)}{n}-\frac{\sum_{i=1}^{n}y_{{}_{i}}(s_{i}=1)}{n}\right|$ (2)
The MLPs were set up the same as in the time experiment. Figure 4 shows the Pareto front, the MO-fANOVA, and the MO-ablation path analysis results for the fairness experiment. While the optimization found several fair configurations, an error remained, with a top accuracy of around 83%.
4.4 Experiment: Energy Consumption
Our last experiment used the Torchvision implementation of ResNets [16] with the first and the last layer set to a size of 3. The models were trained with AdamW and a cosine annealing learning rate scheduler with 200 maximum iterations. They were trained for 50 epochs with early stopping which monitored the test accuracy with a patience of 6 epochs. The CIFAR10 dataset was always normalized with mean and standard deviation. Optionally, the data was augmented with a random crop of size 32 and a random horizontal flip. The objective “inference energy”, so for predicting the test set, was estimated with CodeCarbon [25, 27] [^1] https://github.com/mlco2/codecarbon. The configuration space is displayed in Table 3. Figure LABEL:fig:energy shows the Pareto front, the MO-fANOVA and the MO-ablation path analysis results for the energy experiment.
Table 3: Hyperparameter Space for ResNet
Hyperparameter Belongs to Range Scale Default
layer1 Model 1-30 linear 15
layer2 Model 1-30 linear 15
zero_init_residual Model true or false - true
augment Dataset true or false - false
learning rate Optimizer 0.0001 - 0.1 log 0.01
weight_decay Optimizer 0.00001 - 0.1 log 0.001
eps Optimizer 1e-10-1e-06 log 1e-8
5 Discussion
Our research aimed to investigate whether surrogate-based HPI methods could be converted to meaningfully measure HPI within MOO. We introduced the concept of weighting the objectives based on configurations along the Pareto front and applied it to two HPI methods: fANOVA and ablation path analysis. Subsequently, we conducted three MO HPO experiments with diverse datasets, models, and objectives to assess the usefulness of the two MO-HPI approaches. In the following, we discuss our research questions related to the experimental results from the last section.
5.1 Are the results intuitively correct?
The results show the effectiveness of our method. All three experiments provide plots that make sense from the ML perspective, and thus, we conjecture that the results are intuitively correct. This can be seen in Figure 3 for the ablation and LABEL:fig:energy for fANOVA, where hyperparameters related to network size (such as layer size and number of layers) strongly influence training time and energy consumption, consistent with the understanding that larger neural networks require more energy and time. A similar observation can be made for the learning rate. Its impact on the error is to be expected, given its well-known importance for achieving high accuracy. Additionally, in the second experiment, the learning rate plays a role in fairness (cf. Figure 4). The same is valid for data augmentation in the energy experiment, as is shown in Figure LABEL:fig:energy, it has a high importance and it is known to heavily influence performance [37]. However, some unexpected results, like the high importance of beta_1 in the fairness experiment, are not intuitively explainable and warrant further exploration. We note that recent results showed that improved fairness can be strongly related to hyperparameters [11]. Overall, we therefore claim that the results are intuitively correct, which positively answers our first research question.
5.2 How do the proposed methods compare?
When comparing the two methods, i.e. MO-fANOVA and MO-ablation path analysis, some interesting observations can be made. In the time experiment, the activation function is less important when measured by the ablation path analysis than by fANOVA, but epsilon gains more importance for the error objective in the ablation path analysis. Another surprising observation is that in the fairness experiment, beta_1 is deemed important by both measures but for the fANOVA measure it has a high influence on the error side, and for the ablation path analysis, it has a higher influence on the fairness. Another difference is the high importance of the learning rate in Figure 4 for fANOVA, while there is not much contribution of the learning rate visible for the ablation path analysis. In the energy experiment, in the ablation path analysis data augmentation rated high and the learning rate has minimal relevance in comparison to fANOVA.
The difference in the results of the methods is mostly due to their distinct nature. It is well known that the ablation paths struggle with correctly attributing importance in case of interactions of hyperparameters [14]. Since the value of two or more hyperparameters has to be changed before an effect on the performance can be observed, the importance is only attributed to the last changed hyperparameter. Consequently, the performance differences upon changing the other hyperparameters might be lower than they actually are. In the single objective scenario, this is partially mitigated by also displaying the order of changes which is not possible with our visualisation.
Furthermore, the ablation results strongly depend on the chosen default configuration as a reference point. If the default configuration is already fairly close to one of the extreme ends of the Pareto front, there are little performance gains that can be attributed to the hyperparameters. Some hyperparameters might even seem to be unimportant because they are already set to the optimal value with respect to the dominating objective. Therefore, it is important to always interpret the ablation path in view of the default configuration. Since fANOVA for main effects studies the variance for each hyperparameter independent of the others, it is more stable in that sense. Our method allows us to also study the importance of higher-order interaction effects with fANOVA, as in the single-objective case. To not clutter our plots and results, we have omitted this here.
Lastly, it is important to note that ablation paths are considered a local HPI analysis since their path through the configuration space potentially only covers a small part of it. In contrast, fANOVA covers the entire configuration space and thus provides more high-level insights, which, on the other hand, might not be important for specific performance improvements w.r.t. the default configuration. Overall, it is helpful to use both methods, as they provide diverse insights.
5.3 Can we gain new insights that previous methods could not provide?
One could simply run HPI studies on each objective independently and compare their results. We argue that by considering different tradeoffs of the objectives – as typically done in MOO – we gain valuable insights with our methods that were not possible previously.
Considering the tradeoffs between the objectives, we can make the following observations. In the fANOVA plot of the time experiment (cf. Figure 3), most hyperparameters significantly impacting only one of the objectives intersect at approximately an equal weighting of the objectives. This suggests that the MO-fANOVA in this case contains as much information as calculating fANOVA independently. Nevertheless, even in Figure 3, it is not a trivial linear trend of the importance values of the hyperparameters over different objective tradeoffs. In Figure 4 for the fairness experiment, the intersections are still at around the same position but are not in the center of the plot. For the energy experiment (cf. Figure LABEL:fig:energy), the intersections are even at different x-positions.
In addition, it is evident that certain hyperparameters lose importance rapidly (e.g. the size of Layer 1 in Figure LABEL:fig:energy for the ablation path analysis), while others remain important for longer before they become irrelevant (e.g. beta_1 in Figure 4 for the ablation path analysis). Moreover, some hyperparameters would not be tuned when only looking at the objectives independently. For instance, in the ablation path analysis in Figure 4, the initial learning rate only contributes significantly to the performance within the plot but not on the edges.
Furthermore, it is interesting to observe that certain hyperparameters consistently retain importance, for instance, in Figure 4, for fANOVA both the initial learning rate and the number of neurons remain influential. Conversely, some hyperparameters remain unimportant across all weightings. This consistency provides clarity regarding whether these hyperparameters require tuning. Finally, the stacked visualization makes it possible to easily see the relative contribution of a hyperparameter for different points on the Pareto front which make up the weighting. Therefore, we conclude that there is more information in the plots that considers different objective tradeoffs compared to analyzing the objectives independently.
5.4 Limitations and Future Work
As always being a problem with post-hoc analysis, although our results are reasonable, the absence of ground truth data prevents us from guaranteeing their correctness. Nevertheless, we believe that there is sufficient empirical evidence allowing us to conjecture that our method is correct. Although several approaches for studying (single-objective) results of HPO are already presented [20, 4, 28, 38, 1, 32], a systematic set of criteria as proposed for interpretable ML [29] is still missing in the context of interpretable AutoML.
In our experiments, we have not explored MOO scenarios with more than two objectives. Although our approach can be easily extended to more objectives, plotting the results will be possible for three objectives in a 3D plot, but not beyond this. We note that many-objective optimization (i.e., more than four objectives) could be relevant in practice but leads to having a large fraction of all configurations being on the Pareto front. To the best of our knowledge, there are no reasonable solutions for many-objective HPO to date.
In future work, MO-AutoML could incorporate the proposed method to actively choose the most important hyperparameters to tune. This could be done in the MO setting by calculating the integrals of the MO-HPI over the different tradeoffs. This provides a way to quantify the importance one-dimensionally without having to decide on a tradeoff. This will lead to more efficient AutoML that learns to actively consider the HPI for different tasks, datasets, and models. Moreover, it is already known that post-hoc analysis with partial dependence plots of HPO runs can be skewed because HPO approaches, such as Bayesian optimization [34], tradeoff exploitation and exploration and thus bias the sampled configurations accordingly [28]. So far, it is not known how this affects approaches such as fANOVA and ablation paths, but it is reasonable to assume that they are similarly negatively affected. Therefore, future work has to include how this will affect our approach and new approaches for how to fix skewed results under biased data. In addition, our methods should be integrated into interactive tools, such as DeepCave [33] or IOHprofiler [10] where the weighting could be adjustable by a slider, and other parameters of the method could be adjustable. Another functionality could be enabling users to select configurations on the Pareto front and convert them into objective weightings.
6 Conclusion
In multi-objective hyperparameter optimization, assessing the importance of hyperparameters is challenging due to conflicting objectives. Our proposed method leverages surrogate-based hyperparameter importance measures, specifically fANOVA and ablation paths, in conjunction with a-priori scalarization across various objective tradeoffs. While we focus on those two HPI methods, our approach can also work with other HPI methods, such as investigating the interaction effects of hyperparameters. Empirical evaluations demonstrate the effectiveness of our method. We believe that our proposed approach will enhance the analysis of multi-objective AutoML results.
{ack}
The DFKI Niedersachsen (DFKI NI) is funded in the "zukunft.niedersachsen" by the Lower Saxony Ministry of Science and Culture and the Volkswagen Foundation (funding no. ZN3480). Marius Lindauer was supported by the German Federal Ministry of the Environment, Nature Conservation, Nuclear Safety and Consumer Protection (GreenAutoML4FAS project no. 67KI32007A).